OpenStudy (anonymous):

WILL MEDAL AND FAN The sum of two consecutive odd integers is 124. Find integers. a. n + 3 = 124 b. 2n + 4 = 124 c. 2n + 2 = 124 d. 2n + 1 = 124

OpenStudy (anonymous):

B

OpenStudy (anonymous):

Let \(2n+1\) be the first integer, so the next consecutive odd integer is \(2n+3\). Then \[2n+1+2n+3=124\] Use this to find \(n\), which you can then use to find either integer.

OpenStudy (anonymous):

just go with C because it is always C \[n+n+2=124\\ 2n+2=124\]